# Virtual Target Trajectory Prediction for Stochastic Targets

### v1

@@ 3.5 Outlier Removal
Due to the normalization of the training data, the domain that the CNFs are trained on is limited to $[-1,1]$. Thus, the performance of the CNFs outside this domain can be arbitrarily poor. Furthermore, these samples might even be physically impossible, e.g., if the velocity of the target is assumed to be constant.
To prevent the usage of such outliers, samples outside this range plus a margin of three times the standard deviation of the noise injection are removed from the set of generated samples.
@@ 4 Generation of Virtual Target Trajectories
While the CNFs can be used to predict the distribution of the future positions of the target, the samples generated by the CNFs are not directly usable for downstream tasks, since they are not deterministic trajectories. Besides the development of guidance laws that can handle stochastic trajectories, a more straightforward approach is to generate a set of representative trajectories from the samples, which can be used as a drop-in replacement for deterministic trajectory predictions. To this end, a time series k-means clustering algorithm is used to cluster the samples into a set of representative trajectories. This approach, i.e., the combination of the predictions for multiple targets and the clustering of the samples, as well as the renormalization of the cluster means, is described in the following sections.
@@ 4.1 Combination of Targets
For each of the $n_{\mathrm{r}}$ real targets $i=1,\ldots,n_{\mathrm{r}}$, the respective learned CNFs model is used to generate $j=1,\ldots,n_{\mathrm{s}}$ samples $\mathbf{x}_{i,j,k}$ of the future positions of the target $i$ for each time step $k=1,\ldots,n_{t}$. First, $n_{\mathrm{r}}\cdot n_{\mathrm{s}}$ samples $\mathbf{z}_{i,j}$ are drawn from the base distribution, which is a Gaussian distribution with zero mean and unit variance of dimension $d$, i.e.
@@
$\mathbf{x}_{i,j,k}$ consists of $d$ dimensions, where $d$ is the dimension of the position of the target which can be 2 or 3 depending on the scenario. If the targets are of the same type, i.e., their dynamics and their parameters $\mathbf{\psi}$ are identical, the same generated samples can be used for all the targets to save computation time. Otherwise, a separate CNFs model has to be trained for each type of target and then evaluated for each target. Since most target distributions are not dependent on the current position and azimuth, the samples from the CNFs which are generated in a normalized coordinate system with axis limits of $[-1,1]$ can be moved to the normalized current position $\bar{\mathbf{p}}_{i}$ with orientation $\chi_{i}$ of the real target by performing a rotation and a translation as described in Sec. 3.4.
@@ 4.2 Clustering
After generating and then rotating and translating the samples or each real target $i$, the trajectories $\mathbf{y}_{i,j}$ of the samples are clustered to obtain a set of representative trajectories. To this end, a time series k-means clustering algorithm is used, which assigns each sample trajectory to one of $n_{\mathrm{v}}$ clusters, where $n_{\mathrm{v}}$ is the number of virtual targets that can be chosen by the user. The resulting trajectories of the cluster means are then used as the virtual targets.
The input for the clustering process is a set $Y$ comprised of $n_{\mathrm{r}}\cdot n_{\mathrm{s}}$ flattened trajectories $\mathbf{y}_{i,j}$ of size ($n_{t}\cdot d\times 1$), which are generated as follows:
@@
With this flattened input, a k-means clustering algorithm as described in [19] can be applied to obtain the cluster means, which are then used as the virtual targets.
Figure 6 illustrates the result of the k-means clustering for two dimensions. This only serves as a visualization of the clustering process and does not represent the actual dimensions of the input data, since the input data, i.e., the flattened trajectories, are of size $n_{t}\cdot d$ instead of two dimensions.
Figure 6: Illustration of the resulf of k-means clustering for two dimensions.
By choosing a desired $n_{\mathrm{v}}$, the number of virtual targets, the algorithm assigns each trajectory of the $n_{\mathrm{r}}\cdot n_{\mathrm{s}}$ samples to one of the $n_{\mathrm{v}}$ clusters $C_{i}$ such that the sum of the squared Euclidean distances between the samples and the cluster means is minimized:
$\mathbf{c}_{1}^{*},\ldots,\mathbf{c}_{n_{\mathrm{v}}}^{*}=\mathop{\mathrm{argmin}}_{\mathbf{c}_{1},\ldots,\mathbf{c}_{n_{\mathrm{v}}}}$ $\sum_{i=1}^{n_{\mathrm{v}}}\sum_{\mathbf{y}\in C_{i}}||\mathbf{y}-\mathbf{c}_{i}||^{2}$ (13)
@@
The result of the clustering process is a set of $n_{\mathrm{v}}$ optimal cluster means $\mathbf{c}_{1}^{*},\dots,\mathbf{c}_{n_{\mathrm{v}}}^{*}$, each consisting of dimensions $(n_{t}\cdot d,1)$, which can then be reshaped to $n_{\mathrm{v}}$ trajectories with dimensions $(n_{t},d)$ each:
$\mathbf{c}_{i}^{*}=\begin{bmatrix}\mathbf{c}_{i,1}^{*}&\ldots&\mathbf{c}_{i,d}^{*}\\ \mathbf{c}_{i,d+1}^{*}&\ldots&\mathbf{c}_{i,2d}^{*}\\ \vdots&\ddots&\vdots\\ \mathbf{c}_{i,(n_{t}-1)\cdot d+1}^{*}&\ldots&\mathbf{c}_{i,n_{t}\cdot d}^{*}\end{bmatrix}$ (14)
$\phi_{i}(k\cdot\Delta t)=\left(\mathbf{c}_{i,(k-1)\cdot d+1}^{*},\ldots,\mathbf{c}_{i,k\cdot d}^{*}\right)$ (15)
@@ 4.3 Renormalization
As mentioned in Sec. 3.3, the samples are generated in a normalized coordinate system with axis limits of $[-1,1]$. To obtain the real-world positions and times of the virtual targets, the cluster means have to be renormalized to the original coordinate system, with the inverse of the normalization process mentioned in Sec. 3.5. After applying the renormalization, the trajectories of the virtual targets are obtained, which can be used for further analysis and downstream tasks.
@@ 5 Data Generation
Before training the CNFs model, data has to be generated to train the model. This data consists of trajectories of simulated targets performing random maneuvers generated with a Monte Carlo simulation. If real data is available, it can be used to train the model instead of generating synthetic data.
@@ 5.1 Simple Target
As a practical illustration of the problem described in Sec. 2, consider a target flying in the horizontal plane, i.e., $\mathbf{x}\in\mathbb{R}^{2}$, with a constant velocity and performing random maneuvers. This scenario serves as an instructive example of a stochastically moving target. Three different types of maneuvers are assumed: left turn, right turn, and straight flight. At the beginning of a trajectory (at the origin, flying northbound), the type and duration of the maneuver and the radius (bounded by the lateral acceleration) of the turn are randomly chosen from a uniform distribution with parameters depicted in Table 1. After the maneuver is completed, new maneuvers are randomly chosen until the total duration of the trajectory is reached. Since no additional parameters are required for this simple scenario, $\mathbf{\psi}$ is a vector of dimension zero ($\mathbf{\psi}\in\mathbb{R}^{0}$).
Trajectories created with this approach are shown in Fig. 7. Each dot denotes the change of the maneuver. Since the trajectories are simulated with a time discretization of 0.1 s, trajectory data can be obtained and saved for all the simulated time steps.
Figure 10 depicts histograms of samples of the distribution of the target positions for different times. The data was created with a computationally expensive Monte Carlo simulation.
Table 1: Properties of the target maneuvers. Table reproduced from [18].
@@
[…]
Figure 7: Ten randomly generated simple target trajectories with a duration of 100 s. Figure adapted from [18].
@@ 5.2 Complex Target
$\begin{split}\dot{\mathbf{x}}&=\mathbf{v}\\ \dot{\mathbf{v}}&=\mathbf{g}+\mathbf{d}+\mathbf{w}\\ \end{split}$ (16)
@@
$\mathbf{d}$ $=-\frac{1}{2\cdot\mathrm{BC}}\cdot\rho\cdot||\mathbf{v}||\cdot\mathbf{v}$
@@
The drag $\mathbf{d}$ is calculated with the ballistic coefficient $\mathrm{BC}$, which is the ratio of the mass $m$ of the target to the product of the cross-sectional area $A$ and the drag coefficient $C_{D}$ of the target. The goal of the model is to calculate the distribution of the position $\mathbf{x}$ of the ballistic target after a certain time $t$.
Table 2 depicts the parameters of the ballistic trajectories used as training data. The ballistic coefficient of the target is sampled from the uniform distribution $\mathcal{U}(200,800)~{}\mathrm{kg/m}^{2}$ to allow for the prediction of the distribution of the position of the target for any ballistic coefficient in the learned range. This serves as an example of how additional parameters $\mathbf{\psi}$ can be included in the model to predict the distribution of the target’s position for different parameters. Thus, $\mathbf{\psi}$ is a vector of dimension one ($\mathbf{\psi}\in\mathbb{R}^{1}$) in this scenario.
Table 2: Parameters of the ballistic trajectories used as training data. Table adapted from [18].
@@
Figure 14 depicts samples of the distribution of the target positions for different times and $\mathrm{BC}=500~{}\mathrm{kg/m}^{2}$, obtained by Monte Carlo simulation. It can be seen that the distribution of the target positions widens over time, due to the influence of the noise.
@@ 6.1 Stochastic Predictions
In the following, three different scenarios are considered: First, a stochastically moving target is examined in Sec. 6.1.1. Second, the application of the CNFs to predict trajectories of targets that are moving deterministically is presented in Sec. 6.1.2. Finally, the prediction of the position of a ballistic target with stochastic disturbances is examined in Sec. 6.1.3.
For all three scenarios, the same NNs and CNFs architecture are used, which are depicted in Tables 3 and 4.
@@ 6.1.1 Stochastic Maneuvers
After training the CNFs model for 1000 epochs (requiring 81 s of computation time [^5] All computations were performed with a Ryzen 7 6800U processor with 16 GB RAM.) on the generated data described in Sec. 5 and displayed in Fig. 10, the model is evaluated on the test data. The results of the model evaluation are shown in Fig. 10. The figure depicts the absolute frequency from $10^{4}$ samples (as a representation of the PDF) of the position of the target at different times. Figure 10 depicts the learned PDF of the position of the target at different times.
When comparing the results to the test data depicted in Fig. 10, it can be seen that the model is able to predict the position of the target quite well with a test loss of -2.85. Furthermore, the calculation time for the model evaluation is independent of the time $t$ at which the distribution shall be predicted. Compared to the Monte Carlo simulation, the time to create $10^{4}$ samples is only about 0.17 s, compared to 6.64 s required for the Monte Carlo simulation for the a simulated duration of 100 s. Not only is the sampling much faster, the use of CNFs also allows for an exact computation of the PDF, which is not possible with the Monte Carlo trajectory generation method. Overall, we can conclude that CNFs are a suitable tool for modeling the distribution of the target’s position.
Table 3: Parameters of the Neural Network. Table adapted from [18].
@@
Table 4: Parameters of the Conditional Normalizing Flows model. Table adapted from [18].
@@
Figure 8: Histograms of $10^{4}$ samples of the probability density function of the position of the target at different times obtained by Monte Carlo simulation. Figure reproduced from [18].
@@
Figure 9: Samples of the learned probability density function of the position of the target at different times. Figure reproduced from [18].

### v2

@@ 3.5 Outlier Removal
Due to the normalization of the training data, the domain that the CNFs are trained on is limited to [1,1][-1,1]. Thus, the performance of the CNFs outside this domain can be arbitrarily poor. Furthermore, these samples might even be physically impossible, e.g., if the velocity of the target is assumed to be constant. To prevent the usage of such outliers, samples outside this range plus a margin of three times the standard deviation of the noise injection are removed from the set of generated samples.
@@ 4 Generation of Virtual Target Trajectories
While the CNFs can be used to predict the distribution of the future positions of the target, the samples generated by the CNFs are not directly usable for downstream tasks, since they are not deterministic trajectories. Besides the development of guidance laws that can handle stochastic trajectories, a more straightforward approach is to generate a set of representative trajectories from the samples, which can be used as a drop-in replacement for deterministic trajectory predictions. To this end, a time series k-means clustering algorithm is used to cluster the samples into a set of representative trajectories. This approach, i.e., the combination of the predictions for multiple targets and the clustering of the samples, as well as the renormalization of the cluster means, is described in the following sections. Figure 1 illustrates the structure of the prediction framework, which consists of three main steps: sampling from the CNFs, translating and rotating the samples, and clustering the samples into a set of representative trajectories.
@@ 4.1 Combination of Targets
For each of the $n_{\mathrm{r}}$ real targets $i=1,\ldots,n_{\mathrm{r}}$, the respective learned CNFs model is used to generate $j=1,\ldots,n_{\mathrm{s}}$ samples $\mathbf{x}_{i,j,k}$ of the future positions of the target $i$ for each time step $k=1,\ldots,n_{t}$. First, $n_{\mathrm{r}}n_{\mathrm{s}}$ samples $\mathbf{z}_{i,j}$ are drawn from the base distribution, which is a Gaussian distribution with zero mean and unit variance of dimension $d$, i.e.
@@
$\mathbf{x}_{i,j,k}$ consists of $d$ dimensions, where $d$ is the dimension of the position of the target which can be 2 or 3 depending on the scenario. If the targets are of the same type, i.e., their dynamics and their parameters $\mathbf{\psi}$ are identical, the same generated samples can be used for all the targets to save computation time. Otherwise, a separate CNFs model has to be trained for each type of target and then evaluated for each target. Since most target distributions are not dependent on the current position and azimuth, the samples from the CNFs which are generated in a normalized coordinate system with axis limits of [1,1][-1,1] can be moved to the normalized current position $\bar{\mathbf{p}}_{i}$ with orientation $\chi_{i}$ of the real target by performing a rotation and a translation as described in Sec. 3.4.
@@ 4.2 Clustering
After generating and then rotating and translating the samples or each real target $i$, the trajectories $\mathbf{y}_{i,j}$ of the samples are clustered to obtain a set of representative trajectories. To this end, a time series k-means clustering algorithm is used, which assigns each sample trajectory to one of $n_{\mathrm{v}}$ clusters, where $n_{\mathrm{v}}$ is the number of virtual targets that can be chosen by the user. The resulting trajectories of the cluster means are then used as the virtual targets. The input for the clustering process is a set $Y$ comprised of $n_{\mathrm{r}}n_{\mathrm{s}}$ flattened trajectories $\mathbf{y}_{i,j}$ of size ($n_{t}d\times 1$), which are generated as follows:
@@
With this flattened input, a k-means clustering algorithm as described in [21] can be applied to obtain the cluster means, which are then used as the virtual targets. Figure 6 illustrates the result of the k-means clustering for two dimensions. This only serves as a visualization of the clustering process and does not represent the actual dimensions of the input data, since the input data, i.e., the flattened trajectories, are of size $n_{t}d$ instead of two dimensions.
Figure 6: Illustration of the result of k-means clustering for two dimensions.
By choosing a desired $n_{\mathrm{v}}$, the number of virtual targets, the algorithm assigns each trajectory of the $n_{\mathrm{r}}n_{\mathrm{s}}$ samples to one of the $n_{\mathrm{v}}$ clusters $C_{i}$ such that the sum of the squared Euclidean distances between the samples and the cluster means is minimized:
[…]
@@
The result of the clustering process is a set of $n_{\mathrm{v}}$ optimal cluster means $\mathbf{c}_{1}^{*},\dots,\mathbf{c}_{n_{\mathrm{v}}}^{*}$, each consisting of dimensions $(n_{t}d,1)$, which can then be reshaped to $n_{\mathrm{v}}$ trajectories with dimensions $(n_{t},d)$ each:
$\mathbf{c}_{i}^{*}=\begin{bmatrix}\mathbf{c}_{i,1}^{*}&\ldots&\mathbf{c}_{i,d}^{*}\\ \mathbf{c}_{i,d+1}^{*}&\ldots&\mathbf{c}_{i,2d}^{*}\\ \vdots&\ddots&\vdots\\ \mathbf{c}_{i,(n_{t}-1)d+1}^{*}&\ldots&\mathbf{c}_{i,n_{t}d}^{*}\end{bmatrix}$ (14)
$\phi_{i}(k\Delta t)=\left(\mathbf{c}_{i,(k-1)d+1}^{*},\ldots,\mathbf{c}_{i,kd}^{*}\right)$ (15)
@@ 4.3 Renormalization
As mentioned in Sec. 3.3, the samples are generated in a normalized coordinate system with axis limits of [1,1][-1,1]. To obtain the real-world positions and times of the virtual targets, the cluster means have to be renormalized to the original coordinate system, with the inverse of the normalization process mentioned in Sec. 3.5. After applying the renormalization, the trajectories of the virtual targets are obtained, which can be used for further analysis and downstream tasks.
@@ 5 Data Generation
Before training the CNFs model, data has to be obtained to train the model. This data consists of trajectories of the targets. If real data is available, it can be used to train the model instead of generating synthetic data. However, since real trajectory data for missile targets is not available, synthetic data is generated for the two scenarios described in Sec. 5.1 and Sec. 5.2 using a Monte Carlo simulation and randomly generated target maneuvers.
@@ 5.1 Simple Target
As a practical illustration of the problem described in Sec. 2, consider a target flying in the horizontal plane, i.e., $\mathbf{x}\in\mathbb{R}^{2}$, with a constant velocity and performing random maneuvers. This scenario serves as an instructive example of a stochastically moving target. Three different types of maneuvers are assumed: left turn, right turn, and straight flight. At the beginning of a trajectory (at the origin, flying northbound), the type and duration of the maneuver and the radius (bounded by the lateral acceleration) of the turn are randomly chosen from a uniform distribution with parameters depicted in Table 1 (reproduced from [20]). After the maneuver is completed, new maneuvers are randomly chosen until the total duration of the trajectory is reached. Since no additional parameters are required for this simple scenario, $\mathbf{\psi}$ is a vector of dimension zero ($\mathbf{\psi}\in\mathbb{R}^{0}$). This simple maneuvering logic is maneuver-type independent, meaning that the target’s behavior does not depend on its past behavior, which might not be realistic for all targets, but is sufficient for this simple scenario. If this effect is to be considered, the model can be extended to include the past behavior of the target as additional conditioning variables as demonstrated by [17].
Trajectories created with this approach are shown in Fig. 7 (adapted from [20]). Each dot denotes the change of the maneuver. Since the trajectories are simulated with a time discretization of 0.1 s, trajectory data can be obtained and saved for all the simulated time steps.
Figure 8 (reproduced from [20]) depicts histograms of samples of the distribution of the target positions for different times. The data was created with a computationally expensive Monte Carlo simulation.
Table 1: Properties of the target maneuvers (reproduced from Schneider, M., Loureiro, R., Cunis, T., and Fichter, W., “Trajectory Prediction for Missile Targets: A Probabilistic Approach Using Machine Learning,” in CEAS EuroGNC Conference, 2024. Licensed under a Creative Commons Attribution 4.0 International License ($C$-BY 4.0).).
@@
Trajectory duration $100~\text{s}$
Time discretization $0.1~\text{s}$
Target speed $200~\text{m/s}$
Minimum maneuver duration $5~\text{s}$
Maximum maneuver duration $50~\text{s}$
Minimum lateral acceleration $3~\text{m/s}^{2}$
Maximum lateral acceleration $20~\text{m/s}^{2}$
Figure 7: Ten randomly generated simple target trajectories with a duration of 100 s. Figure adapted from Schneider, M., Loureiro, R., Cunis, T., and Fichter, W., “Trajectory Prediction for Missile Targets: A Probabilistic Approach Using Machine Learning,” in CEAS EuroGNC Conference, 2024. Licensed under a Creative Commons Attribution 4.0 International License ($C$-BY 4.0).
(a) 0 s
(b) 45 s
(c) 90 s
Figure 8: Histograms of $10^{4}$ samples of the probability density function of the position of the target at different times obtained by Monte Carlo simulation. Figure reproduced from Schneider, M., Loureiro, R., Cunis, T., and Fichter, W., “Trajectory Prediction for Missile Targets: A Probabilistic Approach Using Machine Learning,” in CEAS EuroGNC Conference, 2024. Licensed under a Creative Commons Attribution 4.0 International License ($C$-BY 4.0).
@@ 5.2 Complex Target
[…]
@@
[…]
@@
The drag $\mathbf{d}$ is calculated with the ballistic coefficient $\xi$, which is the ratio of the mass $m$ of the target to the product of the cross-sectional area $A$ and the drag coefficient $C_{D}$ of the target. The goal of the model is to calculate the distribution of the position $\mathbf{x}$ of the ballistic target after a certain time $t$.
Table 2 (adapted from [20]) depicts the parameters of the ballistic trajectories used as training data. The ballistic coefficient of the target is sampled from the uniform distribution $\mathcal{U}(200,800)~\mathrm{kg/m}^{2}$ to allow for the prediction of the distribution of the position of the target for any ballistic coefficient in the learned range. This serves as an example of how additional parameters $\mathbf{\psi}$ can be included in the model to predict the distribution of the target’s position for different parameters. Thus, $\mathbf{\psi}$ is a vector of dimension one ($\mathbf{\psi}\in\mathbb{R}^{1}$) in this scenario.
Table 2: Parameters of the ballistic trajectories used as training data (adapted from Schneider, M., Loureiro, R., Cunis, T., and Fichter, W., “Trajectory Prediction for Missile Targets: A Probabilistic Approach Using Machine Learning,” in CEAS EuroGNC Conference, 2024. Licensed under a Creative Commons Attribution 4.0 International License ($C$-BY 4.0).).
@@
Figure 9 (reproduced from [20]) depicts samples of the distribution of the target positions for different times and $\xi=500~\mathrm{kg/m}^{2}$, obtained by Monte Carlo simulation. It can be seen that the distribution of the target positions widens over time, due to the influence of the noise.
(a) 0 s
(b) 45 s
(c) 90 s
Figure 9: Monte Carlo simulation: samples of the probability density function of the position of the target with $\xi=500~\frac{\mathrm{kg}}{\mathrm{m}^{2}}$ at different times. Figure reproduced from Schneider, M., Loureiro, R., Cunis, T., and Fichter, W., “Trajectory Prediction for Missile Targets: A Probabilistic Approach Using Machine Learning,” in CEAS EuroGNC Conference, 2024. Licensed under a Creative Commons Attribution 4.0 International License ($C$-BY 4.0).
@@ 6.1 Stochastic Predictions
In the following, three different scenarios are considered: First, a stochastically moving target is examined in Sec. 6.1.1. Second, the application of the CNFs to predict trajectories of targets that are moving deterministically is presented in Sec. 6.1.2. Finally, the prediction of the position of a ballistic target with stochastic disturbances is examined in Sec. 6.1.3. For all three scenarios, the same NNs and CNFs architecture are used, which are depicted in Tables 3 and 4 (both adapted from [20]).
@@ 6.1.1 Stochastic Maneuvers
After training the CNFs model for 1000 epochs (requiring 81 s of computation time [^5] All computations were performed with a Ryzen 7 6800U processor with 16 GB RAM.) on the generated data described in Sec. 5 and displayed in Fig. 8, the model is evaluated on the test data. The results of the model evaluation are shown in Fig. 11 (reproduced from [20]). The figure depicts the absolute frequency from $10^{4}$ samples (as a representation of the PDF) of the position of the target at different times. Figure 11 (reproduced from [20]) depicts the learned PDF of the position of the target at different times.
When comparing the results to the test data depicted in Fig. 8, it can be seen that the model is able to predict the position of the target quite well with a test loss of -2.85. Furthermore, the calculation time for the model evaluation is independent of the time $t$ at which the distribution shall be predicted. Compared to the Monte Carlo simulation, the time to create $10^{4}$ samples is only about 0.17 s, compared to 6.64 s required for the Monte Carlo simulation for a simulated duration of 100 s. Not only is the sampling much faster, the use of CNFs also allows for an exact computation of the PDF, which is not possible with the Monte Carlo trajectory generation method. Overall, we can conclude that CNFs are a suitable tool for modeling the distribution of the target’s position.
Table 3: Parameters of the Neural Network (adapted from Schneider, M., Loureiro, R., Cunis, T., and Fichter, W., “Trajectory Prediction for Missile Targets: A Probabilistic Approach Using Machine Learning,” in CEAS EuroGNC Conference, 2024. Licensed under a Creative Commons Attribution 4.0 International License ($C$-BY 4.0).).
@@
Table 4: Parameters of the Conditional Normalizing Flows model (adapted from Schneider, M., Loureiro, R., Cunis, T., and Fichter, W., “Trajectory Prediction for Missile Targets: A Probabilistic Approach Using Machine Learning,” in CEAS EuroGNC Conference, 2024. Licensed under a Creative Commons Attribution 4.0 International License ($C$-BY 4.0).).
@@
Figure 10: Samples of the learned probability density function of the position of the target at different times. Figure reproduced from Schneider, M., Loureiro, R., Cunis, T., and Fichter, W., “Trajectory Prediction for Missile Targets: A Probabilistic Approach Using Machine Learning,” in CEAS EuroGNC Conference, 2024. Licensed under a Creative Commons Attribution 4.0 International License ($C$-BY 4.0).
@@
Figure 11: Learned probability density function of the position of the target at different times. Figure reproduced from Schneider, M., Loureiro, R., Cunis, T., and Fichter, W., “Trajectory Prediction for Missile Targets: A Probabilistic Approach Using Machine Learning,” in CEAS EuroGNC Conference, 2024. Licensed under a Creative Commons Attribution 4.0 International License ($C$-BY 4.0).
6.1.2 Deterministic Maneuvers
To prove that a CNFs model with the same architecture can also be used for deterministic maneuvers, it is applied to predict the position of the target described in Sec. 5.1 after a deterministic maneuver of infinite duration. The maneuvers tested were flying straight, a left turn, and a right turn, each with a fixed lateral acceleration of $3~\text{m/s}^{2}$. In essence, this means the target can choose one of three different deterministic trajectories. Compared to the previous scenario, the target dynamics have changed, thus new training data is created in a similar way as before, but with the new target dynamics. Figure 13 (reproduced from [20]) depicts the distribution of the target positions at different instances in the test data.